# Understanding the Impact of Confidence in Retrieval Augmented Generation: A Case Study in the Medical Domain

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) complements the knowledge of Large Language Models (LLMs) by leveraging external information to enhance response accuracy for queries. This approach is widely applied in several fields by taking its advantage of injecting the most up-to-date information, and researchers are focusing on understanding and improving this aspect to unlock the full potential of RAG in such high-stakes applications. However, despite the potential of RAG to address these needs, the mechanisms behind the confidence levels of its outputs remain underexplored. Our study focuses on the impact of RAG, specifically examining whether RAG improves the confidence of LLM outputs in the medical domain. We conduct this analysis across various configurations and models. We evaluate confidence by treating the model’s predicted probability as its output and calculating several evaluation metrics which include calibration error method, entropy, best probability, and accuracy. Experimental results across multiple datasets confirmed that certain models possess the capability to judge for themselves whether an inserted document relates to the correct answer. These results suggest that evaluating models based on their output probabilities determine whether they function as generators in the RAG framework. Our approach allows to evaluate whether the models handle retrieved documents.  [^1] The code is available at https://github.com/naist-nlp/CC_RAG.
@@ 1 Introduction
To address this question, we conduct a systematic analysis of multiple tasks and models in the medical domain, exploring diverse scenarios using PubMedQA Jin et al. (2019) and MedMCQA Pal et al. (2022). In particular, we create psuedo-RAG to manipulate document content – such as adding irrelevant documents deliberately or including only those directly related to the answer – to simulate the range of situations RAG encounter.
@@ 4.4 Evaluation Metrics
Here, $\mathcal{X}$ is the set of all candidate answer choices, and $x^{*}$ denotes the choice with the highest log-probability. Each $v_{i}$ represents the model’s logit for candidate $x_{i}$, and the softmax function converts these logits into a probability distribution over all choices. The selected $x^{*}$ corresponds to the most confident prediction the model makes under the given prompt. This Best Probability reflects how strongly the model favors its top prediction, and it serves as an interpretable confidence score in our evaluations.
Adaptive Calibration Error (ACE)
Adaptive Calibration Error (ACE) Nixon et al. (2019) is a metric proposed to address the shortcomings of Expected Calibration Error (ECE) Naeini et al. (2015); Guo et al. (2017), specifically aiming to reduce the risk of bins with a small number of samples.  Proskurina et al. (2024) and  Ulmer et al. (2022) have pointed out that ACE is a more suitable calibration error metric for multi-class classification problems. Based on these findings, we adopt ACE in our evaluation. Table 2 provides a complete list of all combinations and Appendix A.6 the details of evalution metrics.
@@ 6 Analysis & Discussion
Model Behavior When Inserting Answer-Containing Documents
@@
Behavior of Calibration Error

### v3

@@ Abstract
Retrieval Augmented Generation (RAG) complements the knowledge of Large Language Models (LLMs) by leveraging external information to enhance response accuracy for queries. This approach is widely applied in several fields by taking its advantage of injecting the most up-to-date information, and researchers are focusing on understanding and improving this aspect to unlock the full potential of RAG in such high-stakes applications. However, despite the potential of RAG to address these needs, the mechanisms behind the confidence levels of its outputs remain underexplored. Our study focuses on the impact of RAG, specifically examining whether RAG improves the confidence of LLM outputs in the medical domain. We conduct this analysis across various configurations and models. We evaluate confidence by treating the model’s predicted probability as its output and calculating several evaluation metrics which include calibration error method, entropy, the best probability, and accuracy. Experimental results across multiple datasets confirmed that certain models possess the capability to judge for themselves whether an inserted document relates to the correct answer. These results suggest that evaluating models based on their output probabilities determine whether they function as generators in the RAG framework. Our approach allows us to evaluate whether the models handle retrieved documents. [^1] The code is available at https://github.com/naist-nlp/CC_RAG.
@@ 1 Introduction
To address this question, we conduct a systematic analysis of multiple tasks and models in the medical domain, exploring diverse scenarios using PubMedQA Jin et al. (2019) and MedMCQA Pal et al. (2022). In particular, we create pseudo-RAG to manipulate document content – such as adding irrelevant documents deliberately or including only those directly related to the answer – to simulate the range of situations RAG encounter.
@@ 4.4 Evaluation Metrics
Here, $\mathcal{X}$ is the set of all candidate answer choices, and $x^{*}$ denotes the choice with the highest log-probability. Each $v_{i}$ represents the model’s logit for the candidate $x_{i}$, and the softmax function converts these logits into a probability distribution over all choices. The selected $x^{*}$ corresponds to the most confident prediction the model makes under the given prompt. This Best Probability reflects how strongly the model favors its top prediction, and it serves as an interpretable confidence score in our evaluations.
Adaptive Calibration Error (ACE).
Adaptive Calibration Error (ACE) Nixon et al. (2019) is a metric proposed to address the shortcomings of Expected Calibration Error (ECE) Naeini et al. (2015); Guo et al. (2017), specifically aiming to reduce the risk of bins with a small number of samples.  Proskurina et al. (2024) and  Ulmer et al. (2022) have pointed out that ACE is a more suitable calibration error metric for multi-class classification problems. Based on these findings, we adopt ACE in our evaluation. Table 2 provides a complete list of all combinations and Appendix A.6 the details of evaluation metrics.
@@ 6 Analysis & Discussion
Model Behavior When Inserting Answer-Containing Documents.
@@
Behavior of Calibration Error.